Let ? and ? be two parameters. 

Write the equation system
x1 - s = –x3
3x3 + 3x1 = –x2
rx3 + x2 = x1
 
on the form:




b) We look at the equation system above.

i) Calculate it (A) and show what value r must have for the equation system to have
exactly one solution.

ii) Show what values ​​r and s must have for the system of equations to have infinitely many
solutions.

iii) Show what values ​​r and s must have for the equation system to have no solution.

c) Let F, G and H be three n × n matrices. Solve the following equation with respect to X, when it is
known that the matrix F - G2 is invertible.

F(F + X) = G2(X + H)
Answer #1

X-5 2-23, 3x3 +3%,=-2 823 +22=21 rewrite the aquatiay as xit Hz = 5 (1) thu equations Con be 32,7 22 +373 = 0 & 23 733 put inthe system has no solution if (i) the system has intinitely many solutions a Rank(A)= Rambai <3 . Rank(A) = Rank(AIB)=2 it 8+